OpenStudy (anonymous):

What are the solutions of the quadratic equation below? 2x2 - 8x - 42 = 0

OpenStudy (anonymous):

Factor out 2 o the left side, we get 2(x^2 - 4x - 21) = 0 Factor again (the second expression), we have 2(x - 7)(x + 3) = 0 set each factor equal to 0, we get x = 7, -3 Note, the factor of 2 is never 0 so we ignore that factor.